Groomer’s Report

Weather: -12C, no new snow, total snowpack heights around 280cm, significant rain event on Saturday has left many off-trail areas riding a little crusty after the refreeze on Sunday, trail in excellent condition.

Groomed: Morrissey Trail (three passes to help clean up after last week’s downtime and Saturday’s storm), Summit Trail, Doug Dean Trail, parking lot connector trails. Please note: we are making a push to open Coal Creek for the weekend – we have much work to do here and require all sledders and vehicles to STAYY OFF OF COAL CREEK trail. After today, any vehicles using the staging area or parking on the road should expect to be towed or plowed in behind the grooming until trail officially opens on Saturday. Thank you for your cooperation.

Groomer’s Report

Weather: -7C, clear skies, 5cm of new snow in the past 24 hours, 33cm of new snow in the past 7 days, total snowpack heights around 280cm, major rain event yesterday from valley bottom to above cabin elevations. Significant cooling last night during grooming shift with nice dry snow conditions developing on mid to upper mountain.

Groomed: Morrissey Trail, Doug Dean Trail, Summit Trail, McEvoy Trail, Coal Creek down to the Matheson (Coal Creek Staging Area and the lower trail remain closed at this time for logging activities).

Groomer’s Report

Weather: 5C, 10cm new snow in the last 24 hours, total snowpack heights near 280cm, expect spring-like conditions with rain-soaked snow at all elevations and strong winds. Significant rain event today – please read avalanche bulletin at www.avalanche.ca for more information.

Groomer’s Report

Weather: 0C, snowing heavily to valley bottom, 6cm new snow in the past 24 hours, 33cm in the past 7 days, total snowpack heights around 280cm, all trails in excellent condition.

Groomed: Morrissey Snowcat back online! We’ll be returning to our regular grooming schedule this evening (trail has been groomed this week using the smaller 180 to ‘smoosh-the-whoops’). Corbin groomed last night from the staging area to Rain Gauge.

 

Groomer’s Report

Weather: -4C, clear skies, 4cm new snow in the past 24 hours, 37cm in the last 7 days, total snowpack heights around 280cm.

Conditions: See post below.

Groomer’s Report

Weather: 2C, 4cm in the past 24 hours, 47cm in the past 7 days, total snowpack heights around 3oocm.

Groomed: Big Blue Snowcat is still down for unexpected maintenance (see previous groomers reports). Took the 180 out last night to ‘smush some moguls’ on the Morrissey Trail and Doug Dean Trail.

 

Groomer’s Report

Weather: 0C, 12cm in the past 24 hours, 57cm in the past 7 days, snowpack heights hovering near 300cm.

Groomed: Our Morrissey Groomer is still down for the count resulting in no grooming last night. We choose to focus on initiating troubleshooting and repairs instead of taking out the little 180 and knocking down the whoops. The maintenance issue with the cat appears to be within the pump that builds the high-pressure oil that fires the injectors but requires further testing… a challenging feat to diagnose and repair with the snowcat stuck out in the field. We are hoping to confirm this evening and get parts on their way ASAP. The trail is quite bumpy today (last groom Friday night) but with temperatures hovering around zero they are relatively soft and manageable for most riders at this time. We will re-evaluate the trail conditions and may groom with the little 180 Cat tonight; but, our volunteer’s main focus is directed at getting the big Cat back in service as soon as possible. Thank you for your understanding!

Groomer’s Report

Weather: 2C, 11cm new snow in the past 24 hours, 54cm in the past 7 days, total snowpack heights hovering near 300cm.

Groomed: Morrissey, Summit Trail, Doug Dean, Parking Lot Connector Trail, part of the McEvoy. Snowcat had an unexpected stall and would not restart partway through his groom last night. Downpass on Morrissey Trail not complete. Volunteers are headed up later today to troubleshoot and repair. There may be delays on this evening’s scheduled groom. Please check back for more information.

Groomer’s Report

Conditions: 3C, 15cm new snow in the past 24 hours, 51cm in the past 7 days, snowpack heights hovering around 300cm.

Groomed: Morrissey parking lot connector trails, Corbin Trail

Groomer’s Report

Weather: 1C, 9cm new snow in the past 24 hours, 69cm in the past 7 days, total snowpack heights settling out to between 250-300cm.

Groomed: Morrissey Trail, Doug Dean Trail, Summit Trail, East McEvoy, Notch Trail to the top of the second hill.